OpenStudy (anonymous):

An arithmetic student needs at least a 75% average to receive pass-fail credit for the course. If he scored 85%, 75%, and 84% on the first three exams, what is the lowest score he can get on the fourth exam to receive credit for the course?

OpenStudy (anonymous):

set the equation up: (85+75+84+x)/4=.75 (data 1+data+data n)/n = .75=average where n is 4 because we have 4 data entries. Remember when you get an average= add up your data given/total data entries. Does that make sense? or would you like me to expand? Cross multiply to process and solve for x (isolate x on one side all by itself)...

OpenStudy (anonymous):

Can't make it :(

OpenStudy (anonymous):

you can do it a few ways. if you want to see it better do (85x.25)+(84x.25)+(75x.25) that will give you the current grade. Then just figure 75- that total to get the number you need.

OpenStudy (anonymous):

x=56 ??

Directrix (directrix):

I got 56. (85 + 75 + 84 + x)/4 = 75 244 + x = 300 x = 56

OpenStudy (anonymous):

he requires minimum 75% average to pass then total minimum is 75*4 = 300 when we have assumed total marks per course is equal to 100 and he has scored 85+75+84= 244 in first three courses so he requires 300-244 = 56 minimum to pass